OpenStudy (anonymous):

When using the Babylonian method to approximate ^7, if you start with a guess of 2, then what will be your next guess? a. 4 b. 3.5 c. 3 d. 2.75 anyone know?

OpenStudy (anonymous):



OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if the first guess is \(2\) then the second guess is \[\frac{1}{2}\left(2+\frac{7}{2}\right)\]

OpenStudy (misty1212):

a bit of arithmetic is all that is needed to get the answer

OpenStudy (misty1212):

\[\frac{1}{2}\times \frac{4+7}{2}=\frac{11}{4}\]

OpenStudy (misty1212):

i get \(2.75\)

OpenStudy (anonymous):

thanks guys!! @misty1212 @FateRocks8D

OpenStudy (anonymous):

so @misty1212 would u say its c to?

OpenStudy (misty1212):

\[\color\magenta\heartsuit\] i guess i am a guy no

OpenStudy (anonymous):

@misty1212 is right

OpenStudy (misty1212):

C is 3, but D is \(2.75\)

OpenStudy (anonymous):

oh okok! thanks! @misty1212 @FateRocks8D

OpenStudy (perl):

*